the volume of a certain quantity of n2 gas at 20 Celsius and 100.0 kPa is 15.0L. What will be its volume at 27 Celsius at 1.5 atm?


1
Expert's answer
2022-03-17T09:00:03-0400

P1 = 100.0 kPa × (1 atm / 101.325 kPa) = 0.987 atm

V1 = 15.0 L

T1 = 20°C + 273.15 = 293.15 K

P2 = 1.5 atm

V2 = unknown

T2 = 27°C + 273.15 = 300.15


Solution:

Combined Gas Law can be used:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (0.987 atm × 15.0 L × 300.15 K) / (1.5 atm × 293.15 K) = 10.1 L

V2 = 10.1 L


Answer: The volume will be 10.1 liters